OpenStudy (anonymous):

Multiply and simplify. (2x - 1)(6x - 7) A) 12x2 - 20x - 7 B) 12x2 - 2x + 7 C) 12x2 + 20x + 7 D) 12x2 - 20x + 7

OpenStudy (anonymous):

(2x-1)(6x-7) = 2x(6x) + (2x)(-7) + (-1)(6x) + (-1)(-7) = 12x^2-14x-6x+7 = 12x^2-20x+7

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

should be 12x^2 - 20x + 7

OpenStudy (anonymous):

thanks